OpenStudy (goformit100):

5 * n = 25 n * 5 = 25 (some math rules let you move values around in an equation) (n * 5) ÷ 5 = 25 ÷ 5 (you can divide both sides of an equation by the same value) n * (5 ÷ 5) = 25 ÷ 5 (when using multiplication and division, you can regroup values) n * 1 = 5 (a number divided by itself is equal to one) n = 5 (any value multiplied by one is equal to that value) .... is it correct

OpenStudy (goformit100):

?

OpenStudy (konradzuse):

1. 5 * n = n * 5 I believe is the commutative property of multiplication. 2. Not too sure about the 5/5 business, did you learn that in class? 3. (n / 5 ) * 5 = n (5*5) is the associative property of mult. 4. n * 1 I believe is called the identity property, but not too sure. looks good.

OpenStudy (goformit100):

Thanks @KonradZuse ...